A train leaves Toronto for Montréal at the
same time as another train leaves Montréal
for Toronto. The cities are 500 km apart.
The trains pass each other 2 h later. The
train from Montréal is travelling 50 km/h
faster than the one from Toronto. At what
distance away from Toronto do the trains
pass each other?

I tried to solve the equation but i keep on getting the wrong answer.
Thanks in advance.

Answer by ankor@dixie-net.com(22740)

A train leaves Toronto for Montréal at the same time as another train leaves Montréal for Toronto.
The cities are 500 km apart.
The trains pass each other 2 h later.
The train from Montréal is travelling 50 km/h faster than the one from Toronto.
At what distance away from Toronto do the trains pass each other?
:
Find the speed of the trains first
s = speed of the Toronto train
then
(s+50) = speed of the Montreal train
:
Write a distance equation; dist = time * speed
:
2s + 2(s+50) = 500
2s + 2s + 100 = 500
4s = 500 - 100
4s = 400
s = 400/4
s = 100 km/h is the speed of the Toronto train
:
To find how far he's gone
2 * 100 = 200 mi from Toronto he meets the the other train
:
:
Confirms this by finding out how far the Montreal train has traveled
2 * 150 = 300 mi from Montreal, which is 200 mi from Toronto (500-300)